# SolarGPT-QA: A Domain-Adaptive Large Language Model for Educational Question Answering in Space Weather and Heliophysics

### v2

@@ II-A Forecasting of Solar Flares, CMEs, and SEPs
Existing forecasting approaches can be broadly categorized into empirical, physics-based, and data-driven methods [20]. Empirical techniques exploit correlations between historical proxies, such as sunspot magnetic complexity and X-ray flux, and subsequent eruptions [30, 40, 22]. Physics-based models rely on magnetohydrodynamic simulations and coronal magnetic field extrapolations [60, 64], but their operational use is limited by computational cost and incomplete boundary conditions [63]. Although significant progress has been made in understanding solar magnetic field dynamics, a quantitative theory that rigorously connects the accumulation and restructuring of magnetic flux to the onset of extreme solar events remains lacking, leaving the key mechanisms governing flare and coronal mass ejection initiation still not fully understood [1]. This gap highlights the importance of data-driven approaches and machine learning methods for improving the prediction and characterization of extreme solar events.
Data-driven machine learning approaches have shown substantial promise, particularly for flare and SEP prediction. Multivariate time series models capture the temporal evolution of solar active regions and have demonstrated improved predictive skill over traditional threshold-based methods. Notably, Hamdi et al. [23] formulated solar flare prediction as an end-to-end time series classification problem using photospheric magnetic field parameters. Progress in this area has been supported by the development of benchmark datasets. The Space Weather Analytics for Solar Flares (SWAN-SF) multivariate time series dataset integrates magnetic field measurements with flare labels and has become a widely used resource for solar flare prediction research [2]. Machine learning-based models in this domain include classical approaches such as logistic regression [50], support vector machines [7], and decision trees [33], as well as deep learning architectures for temporal sequence modeling, including LSTMs [36], contrastive learning approaches [57], graph-based methods [59], and transformer frameworks [58]. Subsequent work addressed severe class imbalance and data scarcity, demonstrating that preprocessing, sampling strategies, and data augmentation significantly improve robustness for rare, high-impact events [18, 24].
@@ II-B LLMs for Scientific and Technical Domains
LLMs have evolved rapidly in recent years which can achieve state-of-the-art performance across wide range of the tasks and model scales [8, 14, 54, 55, 11, 10, 9]. There has been subsequent work and exploration for scaling behavior of LLM, architectural variations, and expanded application domains, demonstrating the versatility of LLMs in both general-purpose and specialized settings [44, 67, 53, 46, 66]. DAPT and fine-tuning are common techniques to improve the LLM for specific datasets or domains. Similarly, to improve factual grounding retrieval-augmented generation (RAG) has been used widely which integrates external knowledge sources into language models [28, 26, 19]. In parallel, instruction tuning has emerged as a key alignment technique for adapting LLMs to follow user intent and task-specific objectives [62, 38, 51, 13].
@@ III-C Fine-Tuning on QA dataset
The overall training objective is:
@@
where $\lambda$ controls alignment toward educational clarity.
Unlike prior scientific language models optimized for expert-level analysis, SolarGPT-QA is explicitly designed for educational deployment. By combining domain-adaptive pretraining, parameter-efficient optimization, and constrained decoding, the model bridges expert heliophysics knowledge and accessible K–12 instruction.
@@ IV Experiments
IV-A Training Configuration and Hyperparameters
All experiments were conducted using LoRA-based fine-tuning while keeping the base Meta-LLaMA-3-8B parameters frozen. LoRA adapters were applied to the query and value projection matrices of the attention layers. The rank of the low-rank adapters was set to $r=16$, with a scaling factor $\alpha=32$ and dropout probability $p=0.05$.
During both domain-adaptive pretraining and fine-tuning, optimization was performed using the AdamW optimizer. At inference time, text generation employed stochastic decoding with a temperature of $\tau=0.7$, top-$k=50$, and nucleus sampling with top-$p=0.9$. Generation was terminated at sentence boundaries using a custom stopping criterion, and post-processing filtered incomplete sentence fragments to ensure syntactic completeness and readability. All of our code used in this experiment is available on GitHub [^1] https://github.com/hlabcsusu/SolarGPT.
@@ IV-B1 Domain Corpus for DAPT
Papers were filtered using eruption-related keyword co-occurrence such as solar flare, coronal mass ejection (CME), and solar energetic particles (SEP). Figure 3 shows the distribution of eruption-related research articles across five leading heliophysics journals between 2010 and 2024, stratified by individual and combined event types. This analysis identifies 1,570 relevant papers from these venues alone, and extending the selection to all qualifying journals yields an estimated 5,000–7,000 research articles used for domain adaptation.
@@ IV-B2 QA Dataset Construction
The supervised QA dataset was constructed by generating question–answer pairs from the curated heliophysics literature using large language models, followed by pedagogical rewriting to emphasize clarity, storytelling, and age-appropriate explanations suitable for K–12 learners. Questions focus on conceptual understanding rather than mathematical derivations, targeting physical intuition, cause–effect relationships, and real-world impacts of space weather phenomena.
@@ IV-B3 Train–Test Split and Difficulty Levels
The training set consists of QA pairs derived from the same corpus used during domain-adaptive pretraining. In contrast, the test set is constructed from QA pairs extracted from different PDF documents that are not included in the training corpus, ensuring document-level separation and reducing information leakage. To support a fine-grained educational evaluation, the test set is further divided into three levels of difficulty:
@@ IV-C Baseline
We evaluate SolarGPT-QA against several strong general-purpose large language models that are widely used for scientific question answering and education. The baselines include ChatGPT, Grok, Claude, DeepSeek, and Gemini. All baseline models are evaluated under a zero-shot setting using identical prompts and decoding parameters, and without retrieval augmentation, to ensure a fair and controlled comparison.
In addition, we include an instruction-tuned variant of ChatGPT, where the model is prompt-tuned using a small number of representative examples (few-shot setting), in order to compare SolarGPT-QA against a highly optimized general-purpose model with explicit pedagogical alignment.
@@ IV-D Evaluation Method
We use win rate as the primary evaluation metric based on human pairwise comparison. For each test question, a human evaluator compares SolarGPT-QA’s response with a baseline model’s response and assigns one of three outcomes: Win, Tie, or Loss. The win rate is defined as the proportion of questions where SolarGPT-QA is judged to outperform the baseline. Results are reported by aggregating win, tie, and loss rates across all questions, and separately across difficulty levels (i.e., Easy, Medium, Hard).
@@ V-A Discussion
V-B Future work

### v3

@@ II-A Forecasting of Solar Flares, CMEs, and SEPs
Existing forecasting approaches can be broadly categorized into empirical, physics-based, and data-driven methods [22]. Empirical techniques exploit correlations between historical proxies, such as sunspot magnetic complexity and X-ray flux, and subsequent eruptions [32, 42, 24].Physics-based models rely on magnetohydrodynamic simulations and coronal magnetic field extrapolations [63, 67], but their operational use is limited by computational cost and incomplete boundary conditions [66]. Furthermore, although significant progress has been made in understanding solar magnetic field dynamics, a quantitative theory that rigorously connects the accumulation and restructuring of magnetic flux to the onset of extreme solar events remains lacking in physics-driven approaches, leaving the key mechanisms governing flare and coronal mass ejection initiation still not fully understood [1]. This gap highlights the importance of data-driven approaches and machine learning methods for improving the prediction and characterization of extreme solar events.
Data-driven machine learning approaches have shown substantial promise, particularly for flare and SEP prediction. Multivariate time series models capture the temporal evolution of solar active regions and have demonstrated improved predictive skill over traditional threshold-based methods. Notably, Hamdi et al. [25] formulated solar flare prediction as an end-to-end time series classification problem using photospheric magnetic field parameters. Progress in this area has been supported by the development of benchmark datasets. The Space Weather Analytics for Solar Flares (SWAN-SF) multivariate time series dataset integrates magnetic field measurements with flare labels and has become a widely used resource for solar flare prediction research [2]. Machine learning-based models in this domain include classical approaches such as logistic regression [52], support vector machines [7], and decision trees [35], as well as deep learning architectures for temporal sequence modeling, including LSTMs [38], contrastive learning approaches [59, 60], graph-based methods [62], and transformer frameworks [61]. Subsequent work addressed severe class imbalance and data scarcity, demonstrating that preprocessing, sampling strategies, and data augmentation significantly improve robustness for rare, high-impact events [20, 26].
@@ II-B LLMs for Scientific and Technical Domains
LLMs have evolved rapidly in recent years which can achieve state-of-the-art performance across wide range of the tasks and model scales [8, 16, 56, 57, 13, 12, 11, 10, 9]. There has been subsequent work and exploration for scaling behavior of LLM, architectural variations, and expanded application domains, demonstrating the versatility of LLMs in both general-purpose and specialized settings [46, 70, 55, 48, 69]. DAPT and fine-tuning are common techniques to improve the LLM for specific datasets or domains. Similarly, to improve factual grounding retrieval-augmented generation (RAG) has been used widely which integrates external knowledge sources into language models [30, 28, 21]. In parallel, instruction tuning has emerged as a key alignment technique for adapting LLMs to follow user intent and task-specific objectives [65, 40, 53, 15].
@@ III-C Fine-Tuning on QA dataset
The overall training objective of SolarGPT-QA can be summarized as:
@@
where $\mathcal{L}_{\text{DAPT}}$ denotes the domain-adaptive pretraining loss and $\mathcal{L}_{QA}$ denotes the educational question–answer fine-tuning loss, with $\lambda$ controlling the emphasis on pedagogical alignment. In practice, these objectives are optimized sequentially rather than jointly: the model is first adapted to heliophysics literature using $\mathcal{L}_{\text{DAPT}}$, and the resulting DAPT-adapted checkpoint is subsequently fine-tuned using $\mathcal{L}_{QA}$ to produce the final SolarGPT-QA model. The combined objective summarizes the cumulative learning goal across training stages, rather than a single joint optimization.
Unlike prior scientific language models optimized for expert-level analysis, SolarGPT-QA is explicitly designed for educational deployment. By combining domain-adaptive pretraining, parameter-efficient optimization, and pedagogically aligned fine-tuning with constrained decoding, the model bridges expert heliophysics knowledge and accessible K–12 instruction.
III-D Inference-Time Hallucination Mitigation
In addition to domain-adaptive pretraining and educational fine-tuning, SolarGPT-QA applies inference-time decoding constraints to mitigate hallucinated. These constraints regulate generation behavior rather than model parameters, promoting concise, well-formed, and locally grounded responses suitable for educational settings. Controlled stochastic decoding is used to suppress low-probability and repetitive continuations, while generation length and topic drift are limited through a custom sentence-boundary stopping criterion. Post-generation filtering further removes incomplete or syntactically ill-formed sentence fragments, reducing speculative extrapolation beyond the scope of the input question.
@@ IV Experiments
IV-A Training Details and Compute
Domain-adaptive pretraining (DAPT) and QA fine-tuning were performed using LoRA on Meta-LLaMA-3-8B with 8-bit weight loading and FP16 training. LoRA adapters were applied to the attention q_proj and v_proj modules with rank r=16r{=}16, scaling $\alpha{=}32$, and dropout p=0.05p{=}0.05. Training used the AdamW optimizer with a learning rate of $2\times 10^{-4}$, linear warmup of 100 steps, and a maximum sequence length of 512 tokens. The per-device batch size was 1 with gradient accumulation of 8, resulting in an effective batch size of 8 sequences per optimization step. Both DAPT and QA fine-tuning were run for 3 epochs.
All experiments were conducted on a Linux server with dual Intel Xeon Gold 5220R CPUs (24 cores each, 2.20 GHz) and four NVIDIA RTX A5000 GPUs (24 GB VRAM). Individual training runs used a single GPU, with GPU selection controlled at runtime. To support reproducibility and transparent reporting of compute usage, our training scripts automatically log the total number of optimization steps, wall-clock training time, and peak GPU memory consumption to a training_log.json file. The total number of training tokens can be derived from these logs as:
$\text{tokens}\approx(\text{total steps})\times 512\times 8,$
corresponding to the sequence length and effective batch size used during training. All of our code used in this experiment is available on GitHub [^1] https://github.com/hlabcsusu/SolarGPT.
@@ IV-B1 Domain Corpus for DAPT
Papers were filtered using eruption-related keyword co-occurrence such as solar flare, coronal mass ejection (CME), and solar energetic particles (SEP). Figure 3 shows the distribution of eruption-related research articles across five leading heliophysics journals between 2010 and 2024, stratified by individual and combined event types. This analysis identifies 1,570 relevant papers from these venues alone, and extending the selection to all qualifying journals yields an estimated 5,000–7,000 research articles used for domain adaptation. All documents used were obtained from publicly accessible sources, including open-access journal articles and author-provided PDFs available through publisher websites or institutional repositories. No proprietary databases or restricted-access collections were used. The corpus was used solely for internal model training and was not redistributed in raw or processed form, in accordance with publisher terms of use.
@@ IV-B2 QA Dataset Construction
The supervised QA dataset was constructed using a human-in-the-loop approach and consists of approximately 5,000 question–answer pairs derived from curated heliophysics literature. Large language models were used to generate initial QA candidates, guided by expert-authored few-shot exemplars that defined scientific scope and pedagogical intent. All QA pairs were subsequently reviewed, edited, and validated by human domain experts to ensure factual correctness, appropriate difficulty, and alignment with K–12 learning objectives, while filtering stylistic artifacts associated with synthetic generation.
@@ IV-B3 Train–Test Split and Difficulty Levels
The training set consists of 30 QA pairs derived from the same corpus used during domain-adaptive pretraining. In contrast, the test set is constructed from QA pairs extracted from different PDF documents that are not included in the training corpus, ensuring document-level separation and reducing information leakage. To support a fine-grained educational evaluation, the test set is further divided into three levels of difficulty (10 Easy, 15 Medium, 5 Hard):
@@ IV-C Baseline
We evaluate SolarGPT-QA against several strong general-purpose large language models that are widely used for scientific question answering and education. The baselines include ChatGPT (GPT-4o), Grok (Grok-1.5), Claude (Claude 3 Sonnet), DeepSeek (DeepSeek-V2), and Gemini (Gemini 1.5 Pro). All baseline models are evaluated under a zero-shot setting using identical prompts and decoding parameters, and without retrieval augmentation, to ensure a fair and controlled comparison. In addition, we include an instruction-tuned variant of ChatGPT, where the model is prompt-tuned using a small number of representative examples (few-shot setting), in order to compare SolarGPT-QA against a highly optimized general-purpose model with explicit pedagogical alignment.
@@ IV-D Evaluation Method
We conducted a human pairwise evaluation on 30 test questions, stratified by difficulty (10 Easy, 15 Medium, 5 Hard). For each question, model outputs were independently evaluated by two space weather domain experts, resulting in 60 total judgments per model comparison. Annotators were presented with anonymized and randomly ordered responses and followed a standardized evaluation rubric assessing (i) factual correctness, (ii) clarity of explanation, (iii) age-appropriateness, and (iv) overall educational quality. For each question, annotators assigned one of three outcomes: Win, Tie, or Loss relative to the baseline model. Inter-annotator agreement was measured using Cohen’s $\kappa$, yielding $\kappa=0.72$, which indicates substantial agreement. The win rate is defined as the proportion of questions where model is judged to outperform the baseline. Given the limited number of annotators, the results are indicative rather than statistically conclusive.
@@ V-A Discussion
V-B Limitations
Both training and evaluation QA pairs were initially generated using large language models. Although we enforce strict document-level separation between train and test sources and apply expert human review to reduce synthetic artifacts, residual stylistic overlap or generator-specific priors may inflate performance. Some baseline models share vendors with the QA generators, potentially introducing bias due to shared pretraining or alignment conventions. We partially mitigate this by evaluating multiple independent model families under identical zero-shot settings and by using ablation studies to isolate the effects of domain adaptation and educational fine-tuning. Stronger validation using fully human-authored test sets is left to future work.
V-C Future work